# Stopping Rules for SGD via Anytime-Valid Confidence Sequences

### v4

@@ Abstract
Deciding when to stop stochastic gradient descent (SGD) has long remained unresolved in a statistically rigorous sense. While SGD is routinely monitored as it runs, the classical theory of SGD provides guarantees only at pre-specified iteration horizons and offers no valid way to decide, based on the observed trajectory, when further computation is justified. We address this gap by developing anytime-valid confidence sequences for stochastic gradient methods, which remain valid under continuous monitoring and directly induce statistically valid, trajectory-dependent stopping rules: stop as soon as the current upper confidence bound on an appropriate performance measure falls below a user-specified tolerance. The confidence sequences are constructed using nonnegative supermartingales, are time-uniform, and depend only on observable quantities along the SGD trajectory, without requiring prior knowledge of the optimization horizon. In convex optimization, this yields anytime-valid certificates for weighted suboptimality of projected SGD under general stepsize schedules, without assuming smoothness or strong convexity. In nonconvex optimization, it yields time-uniform certificates for weighted first-order stationarity under smoothness assumptions. We further characterize the stopping-time complexity of the resulting stopping rules under standard stepsize schedules. To the best of our knowledge, this is the first framework that provides statistically valid, time-uniform stopping rules for SGD across both convex and nonconvex settings based solely on its observed trajectory.
@@ 1 Introduction
This issue is particularly acute for stochastic gradient descent (SGD), which in practice is almost always monitored as it unfolds. Existing analyses of SGD characterize its behavior by answering a forward-looking question: if the algorithm is run for a prescribed number of iterations, what guarantees can be made about optimality or stationarity? In practice, however, users face the inverse problem. At a given iteration, after observing the realized trajectory of the algorithm, one would like to know whether continuing to iterate is statistically justified. This inversion of perspective lies at the heart of practical optimization, where computational budgets are finite and decisions must be made online.
@@ 1.1 Problem formulation
Assumption 1.2(i) is the standard unbiasedness condition used in stochastic approximation. Moreover, Assumption 1.2(ii) postulates that the gradient noise is conditionally sub-Gaussian given the past. This condition provides quantitative tail control on the one-step fluctuations of the stochastic gradients and is natural in time-uniform concentration arguments; it is satisfied by bounded, Gaussian, and many light-tailed noise models encountered in practice.
@@
Our goal in the nonconvex setting is to construct time-uniform, data-dependent upper confidence bounds on $\bar{G}_{t}$ that remain valid under continuous monitoring and arbitrary data-dependent stopping.
@@ 1.3 Related work
Time-uniform inference and confidence sequences. Time-uniform (also called anytime-valid) inference provides uncertainty bounds that remain valid simultaneously over all times and are therefore robust to continuous monitoring and optional stopping. A foundational tool is Ville’s inequality for nonnegative supermartingales Ville [1939], which converts supermartingale constructions into time-uniform deviation guarantees. Early statistical instances of this idea include classical confidence sequences for basic parameters Darling and Robbins [1967], as well as boundary-crossing results for martingales Robbins and Siegmund [1970]. Modern developments systematize these principles through nonnegative supermartingale constructions and, in particular, mixture and stitching arguments, yielding sharp time-uniform concentration inequalities that adapt to variance-type processes Howard et al. [2021]; see also related time-uniform Chernoff bounds based on exponential supermartingales Howard et al. [2020] and the broader literature on self-normalized processes De la Pena et al. [2009]. More recently, these ideas have been reframed through e-values and game-theoretic statistics, providing a unified perspective on optional-stopping-safe inference and a comprehensive entry point to the literature Ramdas et al. [2023]; Ramdas and Wang [2025]. While this framework has been developed primarily for sequential testing and estimation, it has begun to influence learning and decision-making problems, including anytime-valid inference for bandits and betting-based estimation Waudby-Smith et al. [2024]; Waudby-Smith and Ramdas [2024]. Our work leverages this time-uniform machinery in an optimization context, using supermartingale mixtures to construct performance certificates for stochastic gradient methods that remain valid uniformly over time under arbitrary, data-dependent stopping.
@@ 2.1 Convex case: suboptimality certificates
$\sigma_{t}^{2}:=\sigma^{2}\eta_{t}^{2}Z_{t},\qquad\Sigma_{t}^{2}:=\sum_{s=1}^{t}\sigma_{s}^{2},\qquad\Sigma_{t,\mathrm{eff}}^{2}:=\max\!\left\{\Sigma_{t}^{2},\;2\Big(\log\tfrac{2}{\alpha}+1\Big)\right\}.$ (5)
@@
Thus, even in this benign one-dimensional model, time-uniform validity forces $U_{t}(\alpha)$ to scale no better than order $1/S_{t}$ along this path. This example illustrates that the appearance of the factor $1/S_{t}$ in time-uniform (anytime-valid) upper bounds is essentially necessary. $\clubsuit$
@@ 2.3 Rates under common stepsize choices
An explicit constant in the bound $\mathbb{E}[U_{t}^{\mathrm{obs}}(\alpha)]=O(1/S_{t})$ can be extracted from the proof. Taken together, Proposition 2.8 and Remark 2.3 show that under stochastic approximation stepsizes, anytime-valid certification comes at essentially no rate cost: the certificate decays as $1/S_{t}$, and this time-uniform scaling is unimprovable in general.
Remark 2.9 (Why no $\log\log\mathrm{t}$ term appears?).
A useful consequence of the square-summable stepsize condition is that the iterated-logarithm correction in $U_{t}^{\mathrm{obs}}(\alpha)$ is expressed in terms of the intrinsic variance scale $\Sigma_{t,\mathrm{eff}}^{2}$ rather than the time index $t$. In classical boundary-crossing problems, the predictable variance typically grows linearly with time, leading to $\log\log\mathrm{t}$ corrections Robbins and Siegmund [1970]; Howard et al. [2021]. Here, however, $\sum_{t=1}^{\infty}\eta_{t}^{2}<\infty$ forces $\Sigma_{t,\mathrm{eff}}^{2}$ to remain uniformly bounded, so the variance-based iterated logarithm is itself bounded. This places stochastic approximation in a regime where variance-adaptive confidence sequences can be substantially tighter than bounds written purely as a function of time; see, for example, De la Pena et al. [2009] and the discussion in [Howard et al., 2021, Section 3]. $\clubsuit$
@@
We briefly comment on the dependence on the confidence level $\alpha$, which is suppressed in the rate statements above. For simplicity, suppose that for the timescales of interest the effective variance proxy satisfies $\widetilde{\Sigma}_{t,\mathrm{eff}}^{2}=\sigma^{2}R_{x}^{2}V_{t}$, that is, the maximum in the definition of $\widetilde{\Sigma}_{t,\mathrm{eff}}^{2}$ is attained by the intrinsic variance scale rather than by the lower truncation term $2(\log({2}/{\alpha})+1)$. In this typical regime, the dependence on $\alpha$ enters only through the boundary term $\log(2/\alpha)$. Under stochastic approximation stepsizes, $V_{t}$ is bounded, and the proof yields
@@ 3 Anytime-Valid Stopping Rules
A central motivation for our confidence-sequence approach is to turn SGD into an anytime-valid optimization procedure: instead of committing to a fixed horizon in advance, the user would like to stop the algorithm based on the information actually observed along the trajectory. In practice, most stopping rules for SGD are heuristic (e.g., they monitor validation loss, gradient norms, or empirical convergence patterns) and any available guarantees typically hold only at a pre-specified time or for a fixed, non-adaptive stopping rule. In contrast, a confidence sequence is designed precisely to support optional stopping: it provides a high-probability guarantee that remains valid under arbitrary data-dependent monitoring and stopping strategies.
As in Section 2.3, we present the stopping rules and their analysis only for the convex confidence sequence from Corollary 2.2, to avoid repetition. At the level of the observable confidence sequences, the convex and nonconvex bounds have identical structure and identical dependence on the stepsize schedule, and differ only in the interpretation of the certified performance measure. In particular, all stopping rules, finiteness arguments, and stopping-time complexity bounds stated below carry over verbatim to the nonconvex confidence sequence from Corollary 2.6, with $\varepsilon$-optimality replaced by $\varepsilon$-stationarity, i.e., with $f(\bar{x}_{t})-f(x^{\star})$ replaced by $\min_{1\leq s\leq t}\|\nabla f(x_{s})\|^{2}$.
@@ 3.1 Stopping-time complexity
To illustrate this, under stochastic approximation stepsizes one has a pathwise envelope of the form $U_{t}^{\mathrm{obs}}(\alpha)\leq K/S_{t}$ almost surely for a random but finite $K$. In this regime, the random prefactor $K$ is essentially given by the terminal value of the nonnegative series $\sum_{t\geq 1}\eta_{t}^{2}\|g_{t}\|^{2}$, up to bounded terms (see the proof of Proposition 2.8). Our assumptions ensure that $K<\infty$ almost surely (and in particular $\mathbb{E}[K]<\infty$), but do not control higher moments of $K$. If in addition $S_{t}\asymp t^{1-\gamma}$ for some $\gamma\in(\tfrac{1}{2},1)$, then Lemma 3.4 yields $\tau_{\varepsilon}\lesssim(K/\varepsilon)^{1/(1-\gamma)}$ almost surely. Consequently, an explicit rate for $\mathbb{E}[\tau_{\varepsilon}]$ requires the moment condition $\mathbb{E}[K^{1/(1-\gamma)}]<\infty$ (for instance, $\gamma=3/4$ requires a fourth moment, while $\gamma=0.9$ requires a tenth moment). This dependence reflects the inversion of the polynomial profile $b(t)=1/S_{t}$, rather than any optional-stopping considerations.
Establishing explicit bounds on $\mathbb{E}[\tau_{\varepsilon}]$ in this regime would therefore require additional moment assumptions on the random prefactor $K$, beyond those needed for anytime-valid certification. Since our aim is to characterize stopping behavior under the same minimal conditions used to construct the confidence sequences, we do not pursue this direction further here. We instead focus on the square-root stepsize schedule $\eta_{t}=\eta_{0}\,t^{-1/2}$, which admits explicit expected stopping-time bounds under the existing assumptions and yields faster decay of the certificate in practice.
Before doing so, for completeness, we also note what can be obtained in the stochastic approximation regime with the stronger assumption that the stochastic gradients are uniformly bounded. In that case, the envelope becomes deterministic, and the certified stopping time can be controlled pathwise rather than only in expectation, yielding explicit trajectory-wise bounds of the type one would classically expect under such assumptions.
Remark 3.5 (SA stepsizes and bounded stochastic gradients).
If one strengthens the gradient assumptions by requiring the stochastic gradients to be almost surely bounded, i.e., $\|g_{t}\|\leq G$, then the envelope in Lemma 3.4 can be taken with a deterministic prefactor, yielding a fully explicit almost sure stopping-time bound under classical stochastic approximation stepsizes. Specifically, under bounded stochastic gradients and $\sum_{t=1}^{\infty}\eta_{t}^{2}<\infty$, the observable certificate admits the pathwise bound
$U_{t}^{\mathrm{obs}}(\alpha)\leq\frac{K}{S_{t}}\qquad\text{almost surely},$
where $K<\infty$ is a deterministic constant depending only on the problem parameters $(G,R_{x})$, the confidence level $\alpha$, and the stepsize schedule through $\sum_{t=1}^{\infty}\eta_{t}^{2}$. Consequently,
$\tau_{\varepsilon}\;\leq\;\inf\!\left\{t\geq 1:\ S_{t}\geq\frac{K}{\varepsilon}\right\}\qquad\text{almost surely}.$
If $\eta_{t}=\eta_{0}\,t^{-\gamma}$ with $\gamma\in(1/2,1)$, then Lemma B.5 implies $S_{t}\gtrsim t^{1-\gamma}$ and hence
$\tau_{\varepsilon}=O\!\left(\varepsilon^{-1/(1-\gamma)}\right)\qquad\text{almost surely}.$
In the boundary case $\eta_{t}=\eta_{0}\,t^{-1}$, Lemma B.6 yields $S_{t}\gtrsim\log t$, and the certified stopping time grows as
$\tau_{\varepsilon}=\exp\bigl(O(\varepsilon^{-1})\bigr).$
These results illustrate that, under strong uniform boundedness assumptions, anytime-valid certification can be combined with classical stochastic approximation stepsizes to recover familiar stopping-time scalings. $\clubsuit$
We now specialize to the square-root schedule $\eta_{t}=\eta_{0}\,t^{-1/2}$. In this regime the deterministic components of $U_{t}^{\mathrm{obs}}(\alpha)$ scale as $(1+\log t)/\sqrt{t}$, and the remaining randomness can be captured by a single weighted noise-square martingale. Lemma B.3 provides the moment control needed to combine Lemma 3.4 with an explicit inversion of the profile $b(t)\asymp(1+\log t)/\sqrt{t}$, yielding an explicit stopping-time complexity rate for $\mathbb{E}[\tau_{\varepsilon}]$.
@@
In Theorem 3.6, the $O(\cdot)$ notation hides constants that depend on the confidence level $\alpha$ and the quantities $(\eta_{0},R_{x},G,\sigma,d)$, but are independent of the tolerance $\varepsilon$. Moreover, the dependence on $\alpha$ enters only through the time-uniform confidence boundary (via $\log(2/\alpha)$ and the lower truncation in the effective variance proxy), and can be made explicit by tracing these factors through the proof; for any fixed $\alpha\in(0,2/\mathrm{e})$ the stated $\varepsilon$-rate is unchanged.
@@
Similarly to Remark 3.5, under the stronger assumption that the stochastic gradients are uniformly bounded, $\|g_{t}\|\leq G$ almost surely, the envelope can be chosen deterministically in the square-root regime as well. In this case one obtains a pathwise certified stopping-time bound (rather than only a bound on $\mathbb{E}[\tau_{\varepsilon}]$), yielding $\tau_{\varepsilon}=O\big(\varepsilon^{-2}\log^{2}(1/\varepsilon)\big)$ almost surely, with constants depending on $(\eta_{0},R_{x},G,\sigma,d,\alpha)$ but not on $\varepsilon$.
@@
The relevance of the process $X_{t}$ emerges once we consider its partial sums
@@
By the conditional sub-Gaussian assumption (17), the conditional expectation in the last line is at most $1$. Hence

### v5

# Stopping Rules for Stochastic Gradient Descent via Anytime-Valid Confidence Sequences

@@ Abstract
The problem of stopping stochastic gradient descent (SGD) in an online manner, based solely on the observed trajectory, is a challenging theoretical problem with significant consequences for applications. While SGD is routinely monitored as it runs, the classical theory of SGD provides guarantees only at pre-specified iteration horizons and offers no valid way to decide, based on the observed trajectory, when further computation is justified. We address this longstanding gap by developing anytime-valid confidence sequences for stochastic gradient methods, which remain valid under continuous monitoring and directly induce statistically valid, trajectory-dependent stopping rules: stop as soon as the current upper confidence bound on an appropriate performance measure falls below a user-specified tolerance. The confidence sequences are constructed using nonnegative supermartingales, are time-uniform, and depend only on observable quantities along the SGD trajectory, without requiring prior knowledge of the optimization horizon. In convex optimization, this yields anytime-valid certificates for weighted suboptimality of projected SGD under general stepsize schedules, without assuming smoothness or strong convexity. In nonconvex optimization, it yields time-uniform certificates for weighted first-order stationarity under smoothness assumptions. We further characterize the stopping-time complexity of the resulting stopping rules under standard stepsize schedules. To the best of our knowledge, this is the first framework that provides statistically valid, time-uniform stopping rules for SGD across both convex and nonconvex settings based solely on its observed trajectory.
@@ 1 Introduction
This issue is particularly acute for stochastic gradient descent (SGD), which in practice is almost always monitored as it unfolds. Existing analyses of SGD characterize its behavior by answering a forward-looking question: if the algorithm is run for a prescribed number of iterations, what guarantees can be made about optimality or stationarity? In practice, however, users face the inverse problem. At a given iteration, after observing the realized trajectory of the algorithm, one would like to know whether continuing to iterate is justified statistically. This inversion of perspective lies at the heart of practical optimization, where computational budgets are finite and decisions must be made online.
@@ 1.1 Problem formulation
Assumption 1.2(i) is the standard unbiasedness condition used in stochastic approximation. Moreover, Assumption 1.2(ii) postulates that the gradient noise is conditionally sub-Gaussian given the past. This condition provides quantitative tail control on the one-step fluctuations of the stochastic gradients and is natural in time-uniform concentration arguments; it is satisfied by bounded, Gaussian, and many light-tailed noise models encountered in practice. In particular, if $\|g_{t}\|\leq G$ almost surely, then (2) holds with $\sigma^{2}:=4G^{2}$.
@@
We note that $\bar{G}_{t}$ is an operational stationarity metric: while small gradient norms certify approximate first-order stationarity, they do not in general provide a calibrated notion of distance to a particular stationary point without additional structure. Our goal in the nonconvex setting is to construct time-uniform, data-dependent upper confidence bounds on $\bar{G}_{t}$ that remain valid under continuous monitoring and arbitrary data-dependent stopping.
@@ 1.3 Related work
Time-uniform inference and confidence sequences. Time-uniform (also called anytime-valid) inference provides uncertainty bounds that remain valid simultaneously over all times and are therefore robust to continuous monitoring and optional stopping. A foundational tool is Ville’s inequality for nonnegative supermartingales Ville [1939], which converts supermartingale constructions into time-uniform deviation guarantees. Early statistical instances of this idea include classical confidence sequences for basic parameters Darling and Robbins [1967], as well as boundary-crossing results for martingales Robbins and Siegmund [1970]. Modern developments systematize these principles through nonnegative supermartingale constructions and, in particular, mixture and stitching arguments, yielding sharp time-uniform concentration inequalities that adapt to variance-type processes Howard et al. [2021]; see also related time-uniform Chernoff bounds based on exponential supermartingales Howard et al. [2020] and the broader literature on self-normalized processes De la Pena et al. [2009]. More recently, these ideas have been reframed through e-values and game-theoretic statistics, providing a unified perspective on optional-stopping-safe inference and a comprehensive entry point to the literature Ramdas et al. [2023]; Ramdas and Wang [2025]. While this framework has been developed primarily for sequential testing and estimation, it has begun to influence learning and decision-making problems, including anytime-valid inference for bandits and betting-based estimation Waudby-Smith et al. [2024]; Waudby-Smith and Ramdas [2024]. Complementing time-uniform validity, expected stopping-time guarantees have also been studied in several strands of the sequential-testing literature. Classical results give bounds on the expected stopping time of the sequential probability ratio test Wald [1992]. More recent work studies expected rejection times for bounded-mean hypothesis testing Chugg et al. [2023]; Chen and Wang [2025], as well as for two-sample and independence testing Shekhar and Ramdas [2023]. Expected rejection-time guarantees have also been developed for sequential hypothesis testing over general classes of composite testing problems Waudby-Smith et al. [2025]. Finally, expected stopping-time guarantees are central in best-arm identification in multi-armed bandits, where sequential tests aim to identify an optimal arm as quickly as possible Garivier and Kaufmann [2016]; Kaufmann et al. [2016]; Kaufmann and Koolen [2021]; Agrawal et al. [2020, 2021]. Our work builds on this time-uniform machinery in an optimization context, using supermartingale mixtures to build SGD performance certificates that are valid under arbitrary, data-dependent stopping and that can also be paired with expected stopping-time analyses under standard stepsize regimes.
@@ 2.1 Convex case: suboptimality certificates
$\sigma_{t}^{2}:=\sigma^{2}\eta_{t}^{2}Z_{t},\qquad\Sigma_{t}^{2}:=\sum_{s=1}^{t}\sigma_{s}^{2},\qquad\Sigma_{t,\mathrm{eff}}^{2}:=\max\!\left\{\Sigma_{t}^{2},\;2\Big(\log\tfrac{2}{\alpha}+1\Big)\right\},$ (5)
@@
Thus, even in this benign one-dimensional model, time-uniform validity forces $U_{t}(\alpha)$ to scale no better than order $1/S_{t}$ along this path. This example illustrates that the appearance of the factor $1/S_{t}$ in time-uniform (anytime-valid) upper bounds is necessary. $\clubsuit$
@@ 2.3 Rates under common stepsize choices
An explicit constant in the bound $\mathbb{E}[U_{t}^{\mathrm{obs}}(\alpha)]=O(1/S_{t})$ can be extracted from the proof. Taken together, Proposition 2.8 and Remark 2.3 show that under stochastic approximation stepsizes, anytime-valid certification comes at essentially no cost to the rate: the certificate decays as $1/S_{t}$, and this time-uniform scaling is unimprovable in general.
Remark 2.9 (Why no $\log\log\mathrm{t}$ term?).
A useful consequence of the square-summable stepsize condition is that the iterated-logarithm correction in $U_{t}^{\mathrm{obs}}(\alpha)$ is expressed in terms of the intrinsic variance scale $\Sigma_{t,\mathrm{eff}}^{2}$ rather than the time index $t$. In classical boundary-crossing problems, the predictable variance typically grows linearly with time, leading to $\log\log t$ corrections Robbins and Siegmund [1970]; Howard et al. [2021]. Here, however, $\sum_{t=1}^{\infty}\eta_{t}^{2}<\infty$ forces $\Sigma_{t,\mathrm{eff}}^{2}$ to remain uniformly bounded, so the variance-based iterated logarithm is itself bounded. This places stochastic approximation in a regime where variance-adaptive confidence sequences can be substantially tighter than bounds written purely as a function of time; see, for example, De la Pena et al. [2009] and the discussion in [Howard et al., 2021, Section 3]. $\clubsuit$
@@
We briefly comment on the dependence on the confidence level $\alpha$, which is suppressed in the rate statements above. For simplicity, suppose that for the timescales of interest the effective variance proxy satisfies $\widetilde{\Sigma}_{t,\mathrm{eff}}^{2}=\sigma^{2}R_{x}^{2}V_{t}$; that is, the maximum in the definition of $\widetilde{\Sigma}_{t,\mathrm{eff}}^{2}$ is attained by the intrinsic variance scale rather than by the lower truncation term $2(\log({2}/{\alpha})+1)$. In this typical regime, the dependence on $\alpha$ enters only through the boundary term $\log(2/\alpha)$. Under stochastic approximation stepsizes, $V_{t}$ is bounded, and the proof yields
@@ 3 Anytime-Valid Stopping Rules
A central motivation for our confidence-sequence approach is to turn SGD into an anytime-valid optimization procedure: instead of committing to a fixed horizon in advance, the user would like to stop the algorithm based on the information actually observed along the trajectory. In practice, most stopping rules for SGD are heuristic (e.g., they monitor validation loss, gradient norms, or empirical convergence patterns) and existing guarantees typically hold only at a pre-specified time or for a fixed, non-adaptive stopping rule. In contrast, a confidence sequence is designed precisely to support optional stopping: it provides a high-probability guarantee that remains valid under arbitrary data-dependent monitoring and stopping strategies.
As in Section 2.3, we present the stopping rules and their analysis only for the convex confidence sequence from Corollary 2.2, to avoid repetition. At the level of the observable confidence sequences, the convex and nonconvex bounds have identical structure and identical dependence on the stepsize schedule, and differ only in the interpretation of the certified performance measure. In particular, all of the stopping rules, finiteness arguments, and stopping-time complexity bounds stated below carry over verbatim to the nonconvex confidence sequence from Corollary 2.6, with $\varepsilon$-optimality replaced by $\varepsilon$-stationarity; i.e., with $f(\bar{x}_{t})-f(x^{\star})$ replaced by $\min_{1\leq s\leq t}\|\nabla f(x_{s})\|^{2}$.
@@ 3.1 Stopping-time complexity
Polynomial stepsizes $\eta_{t}=\eta_{0}\,t^{-1/2}$. We first specialize to the square-root schedule $\eta_{t}=\eta_{0}\,t^{-1/2}$, which admits a fully explicit expected stopping-time bound under the baseline assumptions. In this regime the deterministic components of $U_{t}^{\mathrm{obs}}(\alpha)$ scale as $(1+\log t)/\sqrt{t}$, and the remaining randomness can be captured by a single weighted noise-square martingale. Lemma B.3 provides the moment control needed to combine Lemma 3.4 with an explicit inversion of the profile $b(t)\asymp(1+\log t)/\sqrt{t}$, yielding an explicit stopping-time complexity rate for $\mathbb{E}[\tau_{\varepsilon}]$.
@@
In Theorem 3.5, the $O(\cdot)$ notation hides constants that depend on the confidence level $\alpha$ and the quantities $(\eta_{0},R_{x},G,\sigma,d)$, but are independent of the tolerance $\varepsilon$. Moreover, the dependence on $\alpha$ enters only through the time-uniform confidence boundary (via $\sqrt{\log(2/\alpha)}$ and the lower truncation in the effective variance proxy), and can be made explicit by tracing these factors through the proof; for any fixed $\alpha\in(0,2/\mathrm{e})$ the stated $\varepsilon$-rate is unchanged.
@@
Before proceeding, we note that under the stronger assumption of uniformly bounded stochastic gradients, $\|g_{t}\|\leq G$ almost surely, the square-root regime admits an envelope with a deterministic prefactor. In that case, one also obtains a pathwise certified stopping-time bound, yielding $\tau_{\varepsilon}=O\big(\varepsilon^{-2}\log^{2}(1/\varepsilon)\big)$ almost surely, with explicit constant depending on $(\eta_{0},R_{x},G,\sigma,d,\alpha)$ but not on $\varepsilon$.
Stochastic approximation stepsizes. We now turn to stochastic approximation (SA) stepsizes. Proposition 2.8 yields a pathwise envelope of the same general form, namely $U_{t}^{\mathrm{obs}}(\alpha)\leq K/S_{t}$ almost surely for a finite nonnegative random variable $K$. For polynomial SA stepsizes $\eta_{t}=\eta_{0}\,t^{-\gamma}$ with $\gamma\in(\tfrac{1}{2},1)$, the cumulative weight satisfies $S_{t}\asymp t^{1-\gamma}$, and hence Lemma 3.4 implies $\tau_{\varepsilon}\lesssim(K/\varepsilon)^{1/(1-\gamma)}$ almost surely. Consequently, obtaining an explicit rate for $\mathbb{E}[\tau_{\varepsilon}]$ reduces to integrability of $K^{p_{\gamma}}$ with exponent $p_{\gamma}:=1/(1-\gamma)$ (e.g., $\gamma=3/4$ requires a fourth moment, while $\gamma=0.9$ requires a tenth moment). This dependence is an intrinsic consequence of inverting the polynomial profile $b(t)=1/S_{t}$, rather than of optional stopping.
Under the conditional sub-Gaussian noise assumption, one may control such moments via the same martingale decomposition strategy that underlies the proof of Theorem 3.5: the cumulative noise-square term is decomposed into a predictable compensator plus a martingale $M_{t}$, and martingale maximal/moment inequalities are then used to bound $\sup_{t\geq 1}|M_{t}|$ in $L^{p_{\gamma}}$. In particular, for each fixed $p\geq 2$ one can obtain $\mathbb{E}[K^{p}]<\infty$ by extending the argument of Lemma B.3 to $L^{p}$. The key point, however, is that these bounds are highly non-uniform in $p$. First, high-order control requires correspondingly high moments of $\|\xi_{t}\|$; under conditional sub-Gaussianity these moments are finite for every fixed order, but their deterministic upper bounds grow with the order. Second, the constants in martingale maximal and moment inequalities (e.g., Burkholder’s square-function inequality [Hall and Heyde, 2014, Theorem 2.10] combined with Doob-type maximal inequalities) also increase rapidly with the exponent; in particular, tracking explicit constants typically yields growth on the order of $p^{\,p}$. As $\gamma\uparrow 1$, the required moment order $p_{\gamma}=1/(1-\gamma)$ diverges, and the implicit constants in the resulting bound for $\mathbb{E}[\tau_{\varepsilon}]$ therefore deteriorate extremely fast. Thus, while for each fixed $\gamma\in(\tfrac{1}{2},1)$ one may in principle recover the natural scaling $\mathbb{E}[\tau_{\varepsilon}]=O\big(\varepsilon^{-1/(1-\gamma)}\big)$, the associated constants become quantitatively fragile.
This behavior contrasts with the square-root schedule analyzed in Theorem 3.5. There, the inversion of the profile $b(t)\asymp(1+\log t)/\sqrt{t}$ yields an upper bound of the schematic form $\mathbb{E}[\tau_{\varepsilon}]\lesssim\varepsilon^{-2}\,\mathbb{E}\big[K^{2}\log^{2}(K/\varepsilon)\big]$ (see the proof for the precise statement), so it suffices to control only a moment slightly above $2$. Indeed, for any $\delta>0$, the condition $\mathbb{E}[K^{2+\delta}]<\infty$ implies $\mathbb{E}\big[K^{2}\log^{2}(K)\big]<\infty$, and thus yields a finite bound on $\mathbb{E}[\tau_{\varepsilon}]$ with constants that remain stable.
Therefore, the baseline conditional sub-Gaussian model is sufficient to ensure almost sure termination of the certified stopping rule, but in the SA regime it can lead to expected stopping-time bounds whose constants drastically deteriorate as the stepsizes move from $\eta_{t}=\eta_{0}\,t^{-1/2}$ toward $\eta_{t}=\eta_{0}\,t^{-1}$. To eliminate this issue, one can impose stronger assumptions on the stochastic gradients: if they are uniformly bounded almost surely, then the envelope in Lemma 3.4 is deterministic. As a result, one obtains pathwise certified stopping-time bounds under standard stochastic approximation stepsizes.
Remark 3.7 (SA stepsizes and bounded stochastic gradients).
Assume $\|g_{t}\|\leq G$ almost surely for all $t\geq 1$ and $\sum_{t=1}^{\infty}\eta_{t}^{2}<\infty$. Then there exists a deterministic constant $K_{\mathrm{det}}<\infty$ (depending only on $(G,R_{x},\alpha)$ and the stepsizes through $V_{\infty}:=\sum_{t=1}^{\infty}\eta_{t}^{2}$) such that, almost surely,
$U_{t}^{\mathrm{obs}}(\alpha)\leq\frac{K_{\mathrm{det}}}{S_{t}}\qquad\text{for all }t\geq 1.$
Consequently, almost surely,
$\tau_{\varepsilon}\leq\inf\Bigl\{t\geq 1:\ S_{t}\geq\frac{K_{\mathrm{det}}}{\varepsilon}\Bigr\}.$
If $\eta_{t}=\eta_{0}\,t^{-\gamma}$ with $\gamma\in(1/2,1)$, then Lemma B.5 yields $S_{t}\geq\eta_{0}(t^{1-\gamma}-1)/(1-\gamma)$ for all $t\geq 1$, and hence
$\tau_{\varepsilon}\leq\left\lceil\left(1+\frac{(1-\gamma)K_{\mathrm{det}}}{\eta_{0}\,\varepsilon}\right)^{\!\frac{1}{1-\gamma}}\right\rceil.$
In the boundary case $\eta_{t}=\eta_{0}\,t^{-1}$, Lemma B.6 gives $S_{t}\geq\eta_{0}\log t$, and therefore
$\tau_{\varepsilon}\leq\left\lceil\exp\!\left(\frac{K_{\mathrm{det}}}{\eta_{0}\,\varepsilon}\right)\right\rceil.$
$\clubsuit$
@@
The drift inequality (16) is a standard one-step Lyapunov property for projected stochastic approximation; see, e.g., Nemirovski et al. [2009]. The relevance of the process $X_{t}$ emerges once we consider its partial sums
@@
By the conditional sub-Gaussian assumption (17), the conditional expectation in the last line is at most one. Hence